OpenStudy (anonymous):

A cube has side length eighteen kilometers. What is its surface area? two hundred sixteen square kilometers four hundred thirty-two square kilometers one thousand nine hundred forty-four square kilometers five thousand eight hundred thirty-two square kilometers

OpenStudy (anonymous):

SA = 6s^2 SA = 6(18)^2 SA = 6(324) SA = 1944 In English, this translates to, "one thousand nine hundred forty-four square kilometers".

OpenStudy (anonymous):

thx